# Self-Organising Memristive Networks as Physical Learning Systems

### v1

@@ 2.2 Plasticity
Plasticity refers to the capacity of a system to modulate its internal states or connectivity in response to external stimuli. In biological neuronal networks, plasticity is fundamental to learning and memory, enabling organisms to retain, discard, and process information in response to new experiences. Such adaptability spans multiple timescales and mechanisms, from transient modulation of synaptic strength (short-term plasticity) to longer-lasting changes that underlie memory formation. Physical systems such as SOMNs can similarly exhibit dynamic reconfiguration in response to stimulation, reflecting not only direct, stimulus-specific conductance changes (analogous to homosynaptic plasticity), but also broader network-level adaptations and structural changes (analogous to heterosynaptic plasticity). These capabilities emerge naturally from the interplay of local transport dynamics and the global architecture of the network.
@@
Following experimental demonstrations of memristive devices as synaptic elements in neuromorphic circuits [85, 86], a wide range of synaptic functionalities has been realised in SOMNs due to their capability to dynamically reconfigure in response to external electrical stimuli, particularly in multiterminal configurations. Experimental results show that the emergent behaviour of NW networks can emulate input-specific changes in synaptic conductance pathways, resembling homosynaptic plasticity. Additionally, multiterminal memristive devices demonstrate that plasticity can also extend to synaptic pathways not directly stimulated, effectively mimicking heterosynaptic plasticity effects [44, 87]. These networks also exhibit short-term plasticity, where synaptic weights—represented by the conductance between neuron-like terminals—depend on recent stimulation history, as evidenced by potentiation–during stimulation followed by spontaneous relaxation after the end of stimulation (see Fig. 2(b)). This effect exploits volatile memristive mechanisms inherent in network elements [44, 88, 31]. Short-term memory effects, experimentally demonstrated to be controllable across multiple timescales [44], enable the emulation of working memory processes observed in biological brains. Working memory includes not only the temporary retention of information but also its active manipulation [88, 89]. Crucially, these short-term memory effects facilitate the physical implementation of reservoir computing paradigms, discussed in sec. 4.1.
@@
Beyond short-term effects, NW networks can also preserve previously established pathways while forming new connections. This capability mirrors metaplasticity in biological systems, where synaptic plasticity depends on the historical context of prior modifications [89]. Furthermore, changes in the NW network’s physical topology, known as structural plasticity, have been observed through electrically induced rupture and electrochemical rewiring of NWs [44, 52]. These long-lasting changes in SOMNs can be exploited to emulate long-term plasticity effects, a crucial aspect for storing knowledge in the system during learning. In this context, using tomographical techniques, it has been experimentally demonstrated that spatially correlated short-term plasticity effects can evolve into long-lasting memory traces resembling biological memory engrams, which represent physicochemical changes underlying stored experiences in the brain [56]. Similarly, short-term and long-term memory functionalities have been reported in NP networks[90]. Collectively, these findings demonstrate the potential of SOMNs as substrates for both processing information and storing knowledge at the material level.
In summary, brain–like plasticity effects exhibited by SOMNs reflects their adaptive dynamics and memory properties. While these properties originate from localised memristive junctions, their dynamic network circuitry results in global reconfigurations and associated memory changes. The following section highlights another example of dynamic reconfiguration exhibited by both biological neuronal networks and SOMNs.
@@ 2.3 Criticality
Critical phenomena play a fundamental role in various physical systems, characterised by divergent spatio-temporal correlations and the breakdown of mean-field approximations. Classic equilibrium examples include the paramagnetic–ferromagnetic phase transition described by the Ising model. Importantly, similar critical behaviours emerge in non-equilibrium biological neuronal networks [91], where they are believed to underpin efficient information processing. Here, we highlight that similar criticality has also been experimentally observed in SOMNs, specifically NW and NP networks [16, 75, 92, 93, 94, 95].
@@
NW networks, which are generally fabricated above the percolation density threshold, exhibit criticality when networks are electrically stimulated just above a critical applied voltage [93, 94]. Avalanche criticality observed in Ag-based NW networks is due to electrical percolation when the network dynamics is dominated by electron tunnelling in the junctions [96, 97]. This occurs when junction tunnel gaps form as Ag filaments evolve. These gaps continuously close and re-open as filaments grow and stochastically decay, driving rapid memristive switching between metastable conductance states. At voltages well above the critical voltage, some memristive junctions can form stable filaments that support ballistic transport. When this occurs, NW networks undergo a discontinuous dynamical phase transition resembling dielectric breakdown (also referred to as a “bolt phase”), characterised by the emergence of large amplitude avalanche events [98, 99, 93, 97]. Recent experiments [95] provide deeper insights into critical phenomena in NW networks. These studies demonstrate the programmable induction of critical states via targeted electrical stimulation. Importantly, spatially distinct regions within NW networks can exhibit and sustain local critical dynamics, facilitating spatially correlated interactions that directly influence computational functionality.
@@
While both NW and NP networks show critical dynamics, their measured exponents deviate from known universality classes, hinting at novel behaviour. This may arise due to higher–order interactions or specialised percolation dynamics not fully captured by theoretical percolation models [100, 101], as previously suggested [102, 58, 103]. In particular, the combination of tunneling transport and memristive switching dynamics may distinguish SOMNs from classical percolation universality classes. Recent studies show that tuning network dynamics, via global feedback [104] or local control of junctions [95], can steer systems toward different universality classes, offering pathways to control and apply criticality in SOMNs.
@@
Avalanche criticality in SOMNs is an important concept, as it is a prominent feature that mirrors experiments performed in neuroscience [105]. It is in fact hypothesised that functioning near this point optimises information processing [77, 78, 91]. SOMNs are controllable and experimentally accessible physical systems, and offer valuable abiotic platforms for rigorously testing this hypothesis through direct implementation and analysis of computational tasks (discussed further in sec. 4).
@@ 3 SOMNs as dynamical complex systems
As highlighted in the previous section, SOMNs are complex and dynamic electrical circuits that produce emergent nonlinear dynamics. Here, we present a theoretical framework that explicitly treats SOMNs as dynamical complex systems subject to the physical constraints of Kirchhoff’s electrical conservation laws. This theoretical framework enables models to be developed to gain deeper insights into how the experimentally observable collective dynamics arise from the memristive network interactions.
In particular, these methods are useful for identifying dynamical transitions and to characterise the dynamical regimes that may enhance learning [106, 107, 108], as elaborated in sec. 4.
@@ 3.1 Theoretical Methods
A practical theoretical framework for modelling SOMNs is the lumped circuit approximation. By representing their conductance dynamics using memristive elements, with voltage drops across the junctions, SOMNs can be modelled effectively as assemblies of interacting subcircuits, each corresponding to an individual memristive junction (neglecting the resistance of individual NWs or NPs). Their interactions are governed by Kirchhoff’s laws, resulting in complex, high-dimensional nonlinear dynamical systems.
A convenient way to mathematically capture the diverse nonlinear dynamics observed experimentally is to model each junction’s conductance as a dynamical variable dependent on its voltage or current history. This approach provides a simplified, circuit-inspired model that captures memory and adaptation using only voltage, current, and time, e.g. quantities that are physical observables. This method applies to both NW [109] and NP [110] networks and can be extended to other systems [111].
@@
where $g\in[0,1]$ represents a normalized internal memory parameter, and $v(t)$ denotes the local instantaneous voltage across the junction. The function $F$ encodes how the conductance changes in response not only to $v(t)$ but also to make explicit time dependence, allowing modelling of more complex memory behaviours. Equation $G(g)$ represents the electrical conductance of the junction in terms of $g$. An equivalent representation can also be made in terms of electrical resistance.
While the lumped-element model introduced above is widely applicable (particularly for NW networks), alternative modeling approaches have been developed for NP networks, where different memristive mechanisms are at play. For instance, a potentiation–depression rate balance equation, originally derived from detailed studies of single memristive junctions [112], has been used to describe dynamic conductance changes at both the junction level [27] and in spatially extended networks via a grid-graph modeling framework [87]. These rate-based models offer a complementary theoretical perspective by naturally capturing key plasticity phenomena, including short-term plasticity, metaplasticity, and structural plasticity, as emergent from cumulative electrical activity. As such, they are especially well-suited for describing the temporal evolution of NP networks, where conductance modulation often involves field-induced atomic rearrangement rather than ionic transport mechanisms typical of NW systems.
The lumped-element approximation simplifies the system by reducing each junction to a single memory variable, omitting other physical degrees of freedom that may affect network dynamics. Small but finite capacitive effects at the junctions, though often excluded from basic memristive models, have been experimentally observed and may become significant under certain conditions [113, 114]. These effects introduce distinct memory-like dynamics and can be examined using techniques such as Nyquist impedance spectroscopy.
From a practical standpoint, Kirchhoff’s laws can be numerically solved for a given network and initial junction states to determine voltages and currents across each element. Integrating the dynamical equations (3) then reveals the time evolution of the conductances. This direct approach is widely used in the literature [75, 110, 27, 93, 74] to assess the effective conductance and functional behaviour of SOMNs.
@@
In equations like (4), nonlinear memory effects emerge from the interplay between network topology and time-dependent junction conductances. Unlike resistive networks with fixed internal states, memristive networks adapt their conductance (or resistance) in response to external inputs. The topology is captured via projector matrices, $\Omega$, which typically appear in the equations when Kirchhoff’s laws are implemented analytically also via matrix inverses of the form $(I\pm\chi\Omega\mathcal{G})^{-1}$, where $\mathcal{G}_{ii}=g_{i}$ is a diagonal matrix [115, 109].
@@
This formalism can be extended to more general network motifs, where edges represent subcircuits with complex behaviour, such as memristive, inductive, or capacitive coupling [116]. Though not explored in detail here, such extensions may be important for accurately modelling phenomena like the experimentally observed capacitive effects in SOMNs [113, 114]. Crucially, the matrix inverse structure in the projector formalism induces nontrivial feedback loops within the network. These loops underpin emergent behaviours such as self-organisation and dynamical phase transitions [109], providing a unified analytical framework for linking local memristive dynamics to rich global behaviour.
@@
Incorporating stochasticity into SOMN dynamics offers a powerful extension of the theoretical framework. Stochasticity is a fundamental characteristic of NP networks [75], particularly in the high-voltage regime where the formation and rupture of conductive filaments leads to continuous redistributions of voltage and current, thereby triggering new switching events. The inherent complexity of these networks [58] further implies that the timing of such events is effectively random. Moreover, the filament formation process itself is intrinsically stochastic [76]. Accordingly, SOMNs can be modelled as stochastic dynamical systems [30], with network conductance described, for example, by an Ornstein–Uhlenbeck process that blends stimulus-driven evolution with intrinsic fluctuations. This stochastic mean-field approach captures features such as conductance noise and jumps, while enabling analysis of the potential landscape and the interplay between noise, steady states, and computational capability. In this way, stochastic models provide deeper insight into collective behaviour that cannot be fully described by deterministic frameworks.
The “solid–liquid brain” metaphor [117] illustrates the adaptability of SOMNs: unlike static resistive networks (“solid brains”), SOMNs act as “liquid brains”, flexibly adapting to stimuli, like biological systems [118, 119]. Moreover, equations of the form (4) incorporate memory and topology while enabling Lyapunov function construction, allowing stability analysis akin to Hopfield networks. These functions reveal attractor dynamics in a potential landscape [120], underpinning computational features such as nonlinear responses, metastability, and configuration exploration, which are key for optimisation tasks.

### v2

@@ 2.2 Plasticity
Plasticity refers to the capacity of a system to modulate its internal states or connectivity in response to external stimuli. In biological neuronal networks, plasticity is fundamental to learning and memory, enabling organisms to retain, discard, and process information in response to new signals. Following experimental demonstrations of memristive devices as synaptic elements in neuromorphic circuits [95, 96], a wide range of synaptic functionalities have been realised in SOMNs due to their ability to dynamically reconfigure in response to external electrical stimuli, particularly in multiterminal configurations.
A brief taxonomy. To facilitate the discussion below, we introduce here the main forms of plasticity typical of biological systems that can be implemented in SOMNs. Homosynaptic plasticity: this is a stimulus-specific modulation of synaptic weight (input-specific “synaptic” strengthening/weakening). Heterosynaptic plasticity: a form of synaptic plasticity where the strength of a synapse changes even though that specific synapse was not directly activated. Short-term plasticity (STP): a temporary change in synaptic strength caused by recent presynaptic activity, producing a fading-memory trace of recent inputs. Metaplasticity: plasticity that depends on the prior history of modifications (past activity changes the future plastic response). Structural plasticity: refers to persistent topological changes (creation/removal of conductive connections) that encode longer-term memory.
Homosynaptic and heterosynaptic plasticity in multiterminal SOMNs. Experimental results show that the emergent behaviour of NW networks can emulate input-specific changes in synaptic conductance pathways, resembling homosynaptic plasticity [51, 97]. Additionally, multiterminal memristive devices demonstrate that plasticity can also extend to synaptic pathways not directly stimulated, effectively mimicking heterosynaptic plasticity effects [51, 98]. Heterosynaptic plasticity effects in SOMNs are related to the modulation of not-directly stimulated conductive pathways due to a network-level redistribution of currents/voltages under Kirchhoff constraints.
@@
Short-term plasticity and fading memory (volatile internal state). These networks also exhibit short-term plasticity, where synaptic weights—represented by the effective conductance between neuron-like terminals—depend on recent stimulation history, as evidenced by potentiation during stimulation followed by spontaneous relaxation after the end of stimulation. This effect is related to volatile memristive mechanisms inherent in network elements [51, 99, 33]. Short-term memory effects, experimentally demonstrated to be controllable across multiple timescales [51], enable the emulation of working memory processes observed in biological brains. Working memory includes not only the temporary retention of information but also its active manipulation [99, 100]. Crucially, these short-term, relaxing conductance traces provide the physical basis of a fading-memory kernel, which facilitates the implementation of reservoir computing paradigms (sec. 4.1).
[…]
@@
Metaplasticity, long-term and structural plasticity (persistent memory). Beyond short-term effects, NW networks can also preserve memory of previously established pathways while forming new connections. This capability mirrors metaplasticity in biological systems, where synaptic plasticity depends on the historical context of prior modifications [100]. Furthermore, changes in the NW network’s physical topology, known as structural plasticity, have been observed through electrically induced rupture and electrochemical rewiring of NWs [51, 57]. These long-lasting changes in SOMNs can be exploited to emulate long-term plasticity effects, a crucial aspect for storing knowledge in the system during learning. In this context, tomographic measurements have experimentally shown that spatiotemporal stimulation patterns in multiterminal memristive NW networks can induce adaptive changes in the effective synaptic conductance of the pathways connecting the stimulated terminals, as illustrated in Fig. 2d. These spatially correlated short-term plasticity effects can further evolve into long-lasting memory traces resembling biological memory engrams, namely the physicochemical modifications thought to underlie stored experiences in the brain [62]. Similarly, short-term and long-term memory functionalities have also been reported in NP networks [106]. Collectively, these findings demonstrate the potential of SOMNs as substrates for both processing and storing information at the material level.
In particular, it is possible to observe that: i) stimulation of a pair of electrodes results in potentiation of the corresponding synaptic pathway, followed by spontaneous relaxation due to volatile effects; ii) temporally correlated inputs produce a gradual increase in the effective synaptic conductance, emulating paired-pulse facilitation (PPF); and iii) stimulation of a pair of electrodes also gives rise to changes in non-directly stimulated synaptic pathways (heterosynaptic plasticity) due to the connectivity of the system.
@@ 2.3 Criticality
Critical phenomena play a fundamental role in various physical systems, characterised by divergent spatio-temporal correlations and the breakdown of mean-field theory (MFT) approximations. Classic equilibrium examples include the paramagnetic–ferromagnetic phase transition described by the Ising model. Importantly, similar critical behaviours emerge in non-equilibrium biological neuronal networks [107], where they are believed to underpin efficient information processing.
@@
NW-based SOMNs, which are generally fabricated above the percolation density threshold, exhibit criticality when networks are electrically stimulated just above a critical applied voltage [108, 109]. Avalanche criticality observed in Ag-based NW networks is due to electrical percolation when the network dynamics is dominated by electron tunnelling in the junctions [110, 111]. This occurs when junction tunnel gaps form as Ag filaments evolve. These gaps continuously close and re-open as filaments grow and stochastically decay, driving rapid memristive switching between metastable conductance states. At voltages well above the critical voltage, some memristive junctions can form stable filaments that support ballistic transport. When this occurs, NW networks undergo a discontinuous dynamical phase transition resembling dielectric breakdown (also referred to as a “bolt phase”), characterised by the emergence of large amplitude avalanche events [112, 113, 108, 111].
The first evidence of avalanche criticality in SOMNs was demonstrated in NP networks [20], which are fabricated at the percolation threshold and thus exhibit scale-free responses to electrical stimulation. As shown in Fig. 3, conductance activity patterns (avalanches of switching events) are self-similar across multiple time scales and can be analysed through the probability density function, the inter-event intervals statistics and autocorrelation function.
Recent experiments [114] provide deeper insights into critical phenomena in NW networks. These studies demonstrate the programmable induction of critical states via targeted electrical stimulation. Importantly, spatially distinct regions within NW networks can exhibit and sustain local critical dynamics, facilitating spatially correlated interactions that directly influence computational functionality.
@@
While both NW and NP networks show critical dynamics, their measured exponents deviate from known universality classes, hinting at novel behaviour. This may arise due to higher–order interactions or specialised percolation dynamics not fully captured by standard theoretical percolation models [115, 116], as previously suggested [117, 67, 101]. In particular, the combination of tunneling transport and memristive switching dynamics may distinguish SOMNs from classical percolation universality classes. Recent studies show that tuning network dynamics, via global feedback [118] or local control of junctions [114], can steer systems toward different universality classes, offering pathways to control and apply criticality in SOMNs.
@@
Avalanche criticality in SOMNs is an important concept, as it is a prominent feature that mirrors experiments performed in neuroscience [119]. It is in fact hypothesised that functioning near this point optimises information processing [88, 89, 107, 120]. SOMNs are controllable and experimentally accessible physical systems, and offer valuable abiotic platforms for rigorously testing this hypothesis through direct implementation and analysis of computational tasks (discussed further in sec. 4).
@@ 3 SOMNs as dynamical complex systems
As highlighted in the previous section, SOMNs are complex and dynamic electrical circuits that produce emergent nonlinear dynamics. Here, we present a theoretical framework that explicitly treats SOMNs as dynamical complex systems subject to the physical constraints of Kirchhoff’s electrical conservation laws. This theoretical framework enables models to be developed to gain deeper insights into how the experimentally observable collective dynamics arise from memristive interactions. In particular, this framework is useful for identifying dynamical transitions and to characterise the dynamical regimes that may enhance learning [121, 122, 123], as elaborated in sec. 4. Sec. 3.1 presents a theoretical framework of SOMNs in terms internal memory parameters, while sec. 3.2 presents a macroscopic mean-field description taking into account Kirchhoff’s laws and its relation to experimentally observed collective properties.
@@ 3.1 Theoretical Methods
A practical theoretical framework for modelling SOMNs is the lumped-circuit approximation (i.e. an approximation of the underlying physics in terms of electrical elements and Kirchhoff’s laws), which mathematically describes SOMN networks through a graph-theoretical approach. By representing their conductance dynamics with memristive elements and voltage drops across the junctions, SOMNs can be effectively modelled as assemblies of interacting subcircuits, each corresponding to an individual memristive junction (neglecting the resistance of individual NWs or NPs). Their interactions are governed by Kirchhoff’s laws and constrained by their underlying circuit topology, resulting in complex, high-dimensional nonlinear dynamical systems.
A convenient way to mathematically capture the diverse nonlinear dynamics observed experimentally is to model each junction’s conductance as a dynamical variable dependent on its voltage or current history. This approach provides a simplified model that captures memory and adaptation using only voltage, current, and time, i.e., physical observables. This method applies to both NW [124] and NP [125] networks and can be extended to other systems [126].
@@
where $g\in[0,1]$ represents a normalized internal memory parameter, and $v(t)$ denotes the local instantaneous voltage across the junction. The function $F$ encodes how the conductance changes in response not only to $v(t)$ but also to make explicit time dependence, allowing modelling of more complex memory behaviours. $G(g)$ represents the electrical conductance of the junction in terms of $g$. An equivalent representation can also be made in terms of electrical resistance.
While the lumped-element model introduced above is widely applicable, alternative modelling approaches have been developed, with different memristive mechanisms. For instance, a potentiation–depression rate balance equation, originally derived from detailed studies of single memristive junctions [127], has been used to describe dynamic conductance changes at both the junction level [30] and in spatially extended NW networks via a grid-graph modelling framework [98]. These rate-based models offer a complementary theoretical perspective by naturally capturing key plasticity phenomena, including short-term plasticity, metaplasticity, and structural plasticity, as emergent from cumulative electrical activity.
The lumped-element approximation simplifies the system by reducing each junction to a single memory variable, omitting other physical degrees of freedom that may affect network dynamics. For instance, small but finite capacitive effects at the junctions, though often excluded from basic memristive models, have been experimentally observed and may become significant under certain conditions [128, 129]. These effects introduce distinct memory-like dynamics and can be examined using techniques such as Nyquist impedance spectroscopy [130, 129].
From a practical standpoint, Kirchhoff’s laws can be numerically solved for a given network and initial junction states to determine voltages and currents across each element. Integrating the dynamical equations (3) then reveals the time evolution of the conductances. This direct approach is widely used in the literature [97, 33, 87, 125, 30, 108, 86] to assess the effective conductance and functional behaviour of SOMNs.
@@
In equations like (4), nonlinear memory effects emerge from the interplay between circuit topology and time-dependent junction conductances. Unlike resistive networks with fixed internal states, memristive networks continuously adapt their conductance (or resistance) in response to external driving. The topology is captured via projector matrices, $\Omega$, which typically appear in the equations when Kirchhoff’s laws are implemented analytically also via matrix inverses of the form $(I\pm\chi\Omega\mathcal{G})^{-1}$, where $\mathcal{G}_{ii}=g_{i}$ is a diagonal matrix [131, 124].
@@
This formalism can be extended to more general network motifs, where edges represent subcircuits with more complex, higher-order behaviour, such as inductive or capacitive coupling, in addition to memristive [132]. Though not explored in detail here, such extensions may be important for accurately modelling phenomena like the experimentally observed capacitive effects in SOMNs [128, 129]. Crucially, the matrix inverse structure in the projector formalism induces nontrivial feedback loops within the network. These loops underpin emergent behaviours such as self-organisation and dynamical phase transitions [124], providing a unified analytical framework for linking local memristive dynamics to rich global behaviour.
@@
Incorporating stochasticity into SOMN dynamics offers a powerful extension of the theoretical framework. Stochasticity is a fundamental characteristic of SOMNs [87, 32], particularly in the high-voltage regime where the formation and rupture of conductive filaments leads to continuous redistributions of voltage and current, thereby triggering new switching events. The inherent complexity of these networks [67] further implies that the timing of such events is effectively random. Moreover, the filament formation process itself is intrinsically stochastic [84]. Accordingly, SOMNs can be modelled as stochastic dynamical systems [32], with network conductance described, for example, by an Ornstein–Uhlenbeck process that blends stimulus-driven evolution with intrinsic fluctuations. This stochastic mean-field approach captures features such as conductance noise and jumps, while enabling analysis of the potential landscape and the interplay between noise, steady states, and computational capability. In this way, stochastic models provide deeper insight into collective behaviour that cannot be fully described by deterministic frameworks.
The “solid–liquid brain” metaphor [133] helps clarify, at an intuitive level, why plasticity is central to SOMNs. Unlike static resistive networks (“solid brains”), whose connectivity and functional response are essentially fixed, SOMNs behave more like “liquid brains”: external stimuli continuously reshape their effective conductance pathways (the internal states), allowing the system to adapt its internal state in a way that is closer to biological substrates [134, 135]. These models enable us to describe and even guide the evolution of useful dynamical states, allowing for learning as a form of guided self-organisation [136]. In this sense, plasticity is not an incidental property, but a defining ingredient of the collective dynamics discussed in this section, because it is precisely the coexistence of memory, reconfigurable connectivity, and Kirchhoff-constrained feedback that gives rise to rich emergent behaviour [126] From a theoretical perspective, equations of the form (4) explicitly combine memory and topology, while also allowing the construction of Lyapunov functions and hence a stability analysis akin to that of Hopfield networks. These functions reveal attractor dynamics in an effective potential landscape [137], showing how stimulus-dependent plastic reconfiguration can generate nonlinear responses, metastability, and exploration of multiple configurations. Such features are directly relevant for optimisation and learning, but they also indicate that SOMNs should be viewed as non-equilibrium complex systems in their own right. For this reason, beyond mean-field and circuit-theoretic approaches, novel theoretical methods drawn from complex systems and disordered systems may prove especially valuable for understanding how plasticity, disorder, and feedback jointly shape their emergent dynamics.
The above exposition provides a phenomenological description of SOMNs via Kirchhoff-constrained dynamical equations for the full ensemble of junction states $\vec{g}$. Many experimentally accessible properties, however, concern macroscopic observables such as the effective conductance $G$ and the emergence of collectively distinct low- and high- conductance regimes under changes in control parameters (e.g. voltage, density, connectivity, noise). This motivates a complementary coarse-graining: in the next subsection, we introduce a mean-field theory (MFT) that reduces the high-dimensional network dynamics to an order-parameter description (e.g., $\langle g\rangle$ or $G$), thereby enabling analytical insight into when conductance transitions occur and where the approximation breaks down near critical regimes.